A student is standing on a very slippery surface that is flat and level (perfectly horizontal). She has a mass of 48.0 kg and a backpack with a mass of 12.0 kg. Consider the system to be her and her backpack.

(a)If she throws the backpack horizontally and to her right with a speed of 2.05 m/s relative to her, what is the backpack's speed (in m/s) measured relative to the centre of mass of the system?

(b)What is her speed (in m/s) measured relative to the centre of mass of the system?

(c)If she throws the backpack straight up, is the total momentum of the system conserved?
Answer #1

a) Let her speed be u and bags speed be v relative to center of mass

by momentum conservation,

48u-12v = 0

also, u+v = 2.05

u = 0.41 m/s, v = 1.64 m/s

backpacks speed relative to center of mass =   1.64 m/s

b) her speed = 0.41 m/s

c) it will not be conserved in vertical direction, but will be conserved in horizontal